Calculate the density , specific weight , and volume of chloride gas at 25 ° C and pressure of 600 000 N / m²abs R=118


1
Expert's answer
2021-01-19T10:44:33-0500

Answer:

Initials used;

p for pressure

γ  for density

R for rate



Working


R= "R0 \over M" = "8312 \over 71" = 117.1 N.m/ (kg.k)


γ = "p \over RT" = "600000 N|M2 \over [(117.1 N.m/ (kg.k))] [273+25]" =17.20kg/m3


with g=9.81 m/s2 , y =γg = 17.20 (9.81)= 168.7 N/m3


D= "1 \over γ" = "1 \over 17.20" =  0.0581 m3/kg


Therefore,


Density is 17.20 kg/m3

Specific weight is 168.7 N/m3

Volume of chloride gas 0.0581 m3/kg